Average Rate of Return Method, Net Present Value Method, and Analysis for a service company

The capital investment committee of Arches Landscaping Company is considering two capital investments. The estimated income from operations and net cash flows from each investment are as follows:

Front-End Loader Greenhouse
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $41,800 $127,000 $88,000 $203,000
2 41,800 127,000 67,000 171,000
3 41,800 127,000 33,000 121,000
4 41,800 127,000 15,000 83,000
5 41,800 127,000 6,000 57,000
Total $209,000 $635,000 $209,000 $635,000

Each project requires an investment of $440,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Front-End Loader %
Greenhouse %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Front-End Loader Greenhouse
Present value of net cash flow $ $
Amount to be invested $ $
Net present value $ $

2. Prepare a brief report for the capital investment committee, advising it on the relative merits of the two investments.

The front-end loader has a   net present value because the cash flows occur   in time compared to the greenhouse. Thus, if only one of the two projects can be accepted, the   would be the more attractive.
Answer #1

Average Rate of Return = Average Net Income/Average Investment

Average Investment = (Beginning Investment + Ending Investment)/2

= (440,000+0)/2

= $220,000

Average rate of return

Front end Loader = 41,800/220,000

= 19%

Greenhouse = 41,800/220,000

= 19%

1b. NPV = Present value of cash inflows – Present value of cash outflows

Front end Loader

Greenhouse

Present Value of net cash flow

127,000*PVAF(10%, 5 years)

= 127,000*3.79 = $481,330

508,730

Amount to be invested

440,000

440,000

Net Present Value

41,330

68,730

Lower net present value of front end loader

Later in time

Greenhouse would be more attractive